# The Cycle Counts of Graphs

### v1

@@ Abstract.
An ear in a graph is a path (with distinct endpoints) of positive length whose internal vertices, if any, have degree 2, and whose endpoints have degree greater than 2. When adding an ear, we allow ourselves to subdivide edges to become the endpoints of the ear. It is known that every inseparable graph is a cycle or obtained from a cycle by successively adding ears, and that when we add an ear to an inseparable graph, the result is always inseparable ([5, Theorem 19], [4, Theorem III.11]). The opposite action is deleting an ear, which involves removing the edges and internal vertices of the ear but not its endpoints. We assume this is only done to leave an inseparable remaining graph.
@@ 2. Proof for numbers beyond 89
Now we return to the proof of the main theorem. It was shown in [1] that every integer other than 2, 4, 5, 7, 8, 9, 12, 13, 14, 16, 18, 19, 22, 23, 26, 27, 29, 31, 33, 35, 38, 39, 42, 43, 46, 50, 52, 54, 60, 65, 68, 72, 77 and 89 counts the number of subtrees of a tree. (Cf. [3, Sequence A184164].) By Theorem 7, every positive integer other than those (and 1) is the number of cycles in a 2-connected graph.
@@ 3. Proof for the missing numbers up to 89
A computer search using NAUTY [2] showed that all numbers from 17 up to 100 are cycle counts. We present graphs with the missing cycle counts 22, 23, 26, 27, 29, 31, 33, 35, 38, 39, 42, 43, 46, 50, 52, 54, 60, 65, 68, 72, 77, and 89 in Figure 7. That concludes the proof of Theorem 1.
@@ 4. Proof of Theorem 2
The construction in Section 2 produces inseparable, planar cubic graphs for cycle counts not listed in [1] ([3, Sequence A184164]), as we noted there. This leaves us with the counts listed in [1] apart from 1, 2, 4, 5, 8, 9, 13, and 16. A computer search using NAUTY [2] found planar cubic graphs for these cases, with up to 14 vertices. An example which requires 14 vertices is shown in Figure 8.

### v2

@@ Abstract.
An ear in a graph is a path (with distinct endpoints) of positive length whose internal vertices, if any, have degree 2, and whose endpoints have degree greater than 2. When adding an ear, we allow ourselves to subdivide edges to become the endpoints of the ear. It is known that every inseparable graph is a cycle or obtained from a cycle by successively adding ears, and that when we add an ear to an inseparable graph, the result is always inseparable ([6, Theorem 19], [5, Theorem III.11]). The opposite action is deleting an ear, which involves removing the edges and internal vertices of the ear but not its endpoints. We assume this is only done to leave an inseparable remaining graph.
@@ 2. Proof for numbers beyond 89
Now we return to the proof of the main theorem. It was shown in [1] that every integer other than 2, 4, 5, 7, 8, 9, 12, 13, 14, 16, 18, 19, 22, 23, 26, 27, 29, 31, 33, 35, 38, 39, 42, 43, 46, 50, 52, 54, 60, 65, 68, 72, 77 and 89 counts the number of subtrees of a tree. (Cf. [4, Sequence A184164].) By Theorem 7, every positive integer other than those (and 1) is the number of cycles in a 2-connected graph.
@@ 3. Proof for the missing numbers up to 89
A computer search using NAUTY [3] showed that all numbers from 17 up to 100 are cycle counts. (For computer-friendly data see [2] under the heading “Graphs with given cycle counts”.) We present graphs with the missing cycle counts 22, 23, 26, 27, 29, 31, 33, 35, 38, 39, 42, 43, 46, 50, 52, 54, 60, 65, 68, 72, 77, and 89 in Figure 7. That concludes the proof of Theorem 1.
@@ 4. Proof of Theorem 2
The construction in Section 2 produces inseparable, planar cubic graphs for cycle counts not listed in [1] ([4, Sequence A184164]), as we noted there. This leaves us with the counts listed in [1] apart from 1, 2, 4, 5, 8, 9, 13, and 16. A computer search using NAUTY [3] found planar cubic graphs for these cases, with up to 14 vertices. An example which requires 14 vertices is shown in Figure 8.
@@ 6. Declarations
Data for inseparable planar graphs of order at most 8 having all cycle counts up to 100 are posted at [2] under the heading “Graphs with given cycle counts”.